SpaceRef

SpaceRef


ESA Venus Express Status Report No. 54 Routine Science Operations

Status Report From: ESA Venus Express Mission
Posted: Monday, November 6, 2006

image

21 Nov 2006 09:24

Report for Period 12 November to 18 November 2006

During the reporting period the mission operations have been conducted according to the plan. Science operations have been resumed nominally after the end of the solar conjunction and the work-around for the TM burst generation in VIRTIS seems to be effective.

On DoY 318 a problem in the ground station has caused the loss of the complete ground station pass. This has had no effect on the operations.

The table below shows a chronology of the main activities in the reporting period:

MET (Day) Date DOY VPER# Main Activity
369 12/11/06 316 205 Routine Mission operations
370 13/11/06 317 206 Routine Mission operations
371 14/11/06 318 207 Routine Mission operations, pass skipped due to GS failure
372 15/11/06 319 208 Routine Mission operations
373 16/11/06 320 209 Routine Mission operations
374 17/11/06 321 210 Routine Mission operations
375 18/11/06 322 211 Routine Mission operations

At the end of the last Cebreros pass in the reporting period (DOY 322, 16:00z) Venus Express was orbiting Venus at 254 million km from the Earth. The one-way signal travel time was 850 seconds.


Payload Activities

ASPERA
The instrument is regularly operated as part of the routine plan.

MAG
The instrument is regularly operated as part of the routine plan.

PFS
The instrument is OFF and is not operated.

SPICAV
The instrument is regularly operated as part of the routine plan.

VeRA
The USO is nominally powered but muted. RSI operations, when conducted, are done with TTC chain 1, meaning without USO.

VIRTIS
The instrument is regularly operated as part of the routine plan.

VMC
The instrument is regularly operated as part of the routine plan.


Mission Planning

Short Term Planning
Operations of STP028 were executed successfully and the planning of STP029 was completed in time.

The science data dump has been delayed by a missed ground station contact on DoY 318 but no data has been lost.

Medium Term Planning
MTP010 planning has been successfully completed.

Long Term Planning
A decision whether to request DSN support for the movie observation, to be performed in April 2007, has not been taken yet.


Future Milestones

The nominal spacecraft configuration and the mission operations are being fully resumed after the superior conjunction phase. The normal pace of operations will now be followed again.

The second Earth occultation season will start in December with additional Radio Science operations with the ESA New Norcia station.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter